# BCR-DRL: Behavior- and Context-aware Reward for Deep Reinforcement Learning in Human-AI Coordination

### v4

@@ Abstract
Deep reinforcement Learning (DRL) offers a powerful framework for training AI agents to coordinate with human partners. However, DRL faces two critical challenges in human-AI coordination (HAIC): sparse rewards and unpredictable human behaviors. These challenges significantly limit DRL to identify effective coordination policies, due to its impaired capability of optimizing exploration and exploitation. To address these limitations, we propose an innovative behavior- and context-aware reward (BCR) for DRL, which optimizes exploration and exploitation by leveraging human behaviors and contextual information in HAIC. Our BCR consists of two components: (i) A novel dual intrinsic rewarding scheme to enhance exploration. This scheme composes an AI self-motivated intrinsic reward and a human-motivated intrinsic reward, which are designed to increase the capture of sparse rewards by a logarithmic-based strategy; and (ii) A new context-aware weighting mechanism for the designed rewards to improve exploitation. This mechanism helps the AI agent prioritize actions that better coordinate with the human partner by utilizing contextual information that can reflect the evolution of learning in HAIC. Extensive simulations in the Overcooked environment demonstrate that our approach can increase the cumulative sparse rewards by approximately $20\%$ and reduce the convergence time by about $67\%$ compared to state-of-the-art baselines.
@@ 1 Introduction
Human-AI coordination (HAIC) has emerged as a critical research area focusing on complicated tasks requiring coordinated behaviors, including the synergy of human intuition and machine autonomy Carroll et al. (2019); Sarkar et al. (2022), thereby pressing the need of an AI agent adaptive to the human partner in addition to the environment. Deep reinforcement learning (DRL) offers a powerful framework for developing such an adaptive AI agent, leveraging its ability to learn optimal policies through interactions with both the human partner and the environment Zhao et al. (2023); Loo et al. (2023). However, employing DRL in HAIC presents two interdependent challenging goals in balancing exploration and exploitation:
@@
i) Enhancing exploration of critical but rare state-action pairs yielding sparse rewards. Rewards are usually sparse in HAIC due to the inherent complexity of coordinated tasks. The state-action pairs associated with sparse rewards in HAIC are critical for effective learning. However, these pairs are rarely encountered during training, since they emerge from coordinated sequences of temporally-extended actions between AI and human agents Goecks et al. (2019)–making them particularly elusive when the AI agent has limited ability to fully interpret the intricate dynamics of unpredictable human behaviors Ghosal et al. (2023); Zhao et al. (2020). These rare interactions represent key moments when AI actions perfectly align with human intentions, leading to successful task completion. Without sufficient exposure to these critical state-action pairs, the AI agent struggles to understand which actions are beneficial, resulting in slow learning and local optima performance. Existing approaches mitigate this issue by augmenting exploration through intrinsic rewards that supplement those sparse rewards obtained extrinsically from the environment Jaques et al. (2019); Haarnoja et al. (2018).
@@
ii) Ensuring effective exploitation of the explored state-action pairs. Although the above sparse rewarding challenge in HAIC can be mitigated by intrinsic rewards, continuously relying on them throughout training can lead to unstable performance Lowe et al. (2017); Hadfield-Menell et al. (2016); Pathak et al. (2017), due to the inherent exploration-exploitation trade-off in DRL Sutton and Barto (2018); Mnih et al. (2016); Rashid et al. (2020). To address this issue, we can leverage the training context in HAIC to optimize the exploration and exploitation. Specifically, the context in HAIC captures the evolving state of coordination between AI and human participants, such as accumulated sparse rewards, rewards instability conditions, and learning progress. This contextual information enables AI agents to adaptively adjust their learning strategy: prioritizing exploration in early stages when critical state-action pairs are scarce, while gradually transitioning to exploitation in later stages to refine the learned coordination patterns.
@@
In this paper, we design an innovative behavior- and context-aware reward (BCR) for DRL, namely BCR-DRL, to address the aforementioned challenges in HAIC. Our BCR (see Figure 1) extends conventional extrinsic rewards with two key components: (i) Dual intrinsic rewards enhance exploration by encouraging both diverse AI actions and distinct behavior from the human partner, and (ii) Context-aware weights that optimize exploitation by dynamically adjusting reward weights based on the training context. Specifically, the first component includes two intrinsic rewards: AI self-motivated rewards that promote action diversity, and human-aware rewards that help understand human intentions through counterfactual reasoning. Both rewards utilize logarithmic terms to increase the likelihood of encountering critical but rare state-action pairs. The second component dynamically balances intrinsic and extrinsic rewards based on coordination effectiveness, measured through task completion rates, human behavior patterns, and agent diversity. Experimental results demonstrate that our approach outperforms state-of-the-art algorithms, achieving higher sparse rewards and faster training convergence, which illustrate an innovative path for effective and seamless HAIC.
@@ 2 Related Work
Intrinsic rewards.
Prior work in multi-agent reinforcement learning has demonstrated the effectiveness of social influence as intrinsic rewards Jaques et al. (2019). This approach uses counterfactual actions of the ego AI agent to encourage behaviors that significantly influence other AI agents’ actions. While effective in pure AI scenarios, this mechanism faces challenges in HAIC where human behavior patterns are inherently less predictable than those of AI agents. This unpredictability can lead to suboptimal policies when the AI agent attempts to influence human actions, caused by inadequate exploration. More recent work has proposed a model-free approach that maximizes AI behavior diversity without explicitly modeling human behavior Zhao et al. (2023). While this approach enables adaptation to general human behaviors, the no free lunch theorems Wolpert and Macready (1997) suggest that optimal performance in specific scenarios requires exploiting knowledge about the particular behavioral patterns of individuals or groups.
We bridge this gap by proposing a dual intrinsic rewarding scheme that combines the benefits of both approaches while addressing their limitations. Our approach composes: (i) an AI self-motivated intrinsic reward that facilitates comprehensive exploration through behavioral diversity, and (ii) a human-motivated intrinsic reward that utilizes counterfactual human actions obtained from the pre-trained human model, enabling the AI agent to better adapt human intentions. This dual approach enables the AI agent to maintain a comprehensive exploration of its action space while developing actions that effectively complement specific human behavioral patterns.
Figure 1: The proposed BCR-DRL for HAIC. The AI agent is trained by our BCR-DRL algorithm, which integrates extrinsic rewards, intrinsic rewards, and context-aware weights. Extrinsic rewards are obtained from the environment, whilst intrinsic rewards are developed based on the behaviors of both the human and AI agents. Context-aware weights are dynamically adjusted according to the training context, such as the episode return–across three domains task performance, AI agent behaviors, and human behaviors.
Critical rare state-action pairs.
Capturing the critical state-action pairs associated with the sparse rewards is crucial for efficient DRL training in complicated scenarios like HAIC. However, these pairs are usually rarely encountered since they are the results of executing a series of temporally-extended actions Goecks et al. (2019); Ghosal et al. (2023); Zhao et al. (2020). Without sufficient exposure of these critical rare state-action pairs, the AI agent struggles to identify effective policies, resulting in slower learning and local optima performance.
@@
Although incorporating intrinsic rewards can encourage the exploration of the AI agent, it introduces the risk of over-exploration and low data sample utilization efficiency. This issue becomes particularly acute when critical state-action pairs, which were initially rare, but being captured more frequently as training progresses. At this point, continuing to emphasize exploration through intrinsic rewards may hinder rather than help the learning process. Therefore, a desirable DRL reward should be context-aware–capable of recognizing and adjusting when to emphasize the exploration by enlarging the intrinsic rewards and when to emphasize the exploitation by strengthening the extrinsic rewards.
@@
The key challenge here lies in dynamically balancing these competing objectives based on the current training context. This bears a similarity to balancing multiple tasks in deep learning. For example, adaptive weighting has been developed to dynamically adjust task weights based on the instabilities caused by the co-existing tasks in deep learning Liu et al. (2019). Inspired by this concept of adaptive weighting, we propose a novel context-aware weighting mechanism for HAIC. Our approach dynamically adjusts the weights between different rewards by monitoring the training context, including the accumulated sparse rewards and instabilities caused by different rewards, aiming at efficiently achieving the global optimal of sparse rewards.
3 Benchmark Human-AI Coordination
Experimental coordination aspects.
@@ 3 Benchmark Human-AI Coordination
We evaluate our approach using Overcooked Carroll et al. (2019); Carroll (2024), a standardized benchmark for HAIC that simulates collaborative cooking tasks requiring coordinated actions between human and AI agents. This environment is particularly suitable for studying HAIC emphasizing sparse rewards, as successful task completion requires executing precise sequences of coordinated actions of both human and AI agents. We conduct experiments across three distinct layouts of the overcooked environment. Each of the three layouts highlights different aspects of coordination between human and AI agents. The Cramped Room layout focuses on coordination between agents located in a shared space, where collision avoidance between human and AI agents is critical; The Asymmetric Advantages layout concentrates on coordination between agents located in distinct areas with varying access to cooking resources, where asymmetric behavior planning is critical; The Coordination Ring layout evaluates coordination between agents located in a small room with a central obstacle, where path planning is critical. Detailed analysis along with the experiments are given in Section 5.
Human behavior alignment.
@@
Current DRL-based AI agents for HAIC fall into two categories: model-free approaches that train AI agents to adapt to general human behaviors without relying on specific human model Zhao et al. (2023), and model-based approaches that enable more personalized adaptation to individual or group-specific behavioral patterns Carroll et al. (2019); Loo et al. (2023); Buçinca et al. (2024). Our work adopts the model-based approach. The human models utilized in this paper are developed by the authors of Carroll et al. (2019), which use behavior cloning to imitate the behaviors of real human players in the Overcooked environment. We aim to align these human models across different layouts using our BCR-DRL algorithm designed for the AI agent.
@@
The selection of the model-based approach ensures reliable validation of our BCR-DRL algorithm, particularly its human-motivated intrinsic reward component which builds upon accurate modeling of human behavior. The three layouts are selected based on their demonstrated reliability in previous work where trained human models achieved performance comparable to average human players Carroll et al. (2019).

### v5

@@ Abstract
[…]
123
@@ 1 Introduction
Human-AI coordination (HAIC) has emerged as a critical research area focusing on complicated tasks requiring coordinated behaviors, including the synergy of human intuition and machine autonomy [1, 2]. For instance, in the Overcooked environment, an AI agent must coordinate with a human partner to prepare meals by dividing tasks such as chopping ingredients and delivering dishes, requiring adaptive responses to human actions (detailed in Section 3.2). This pressing need demands an AI agent that is adaptive to the human partner in addition to the environment. Deep reinforcement learning (DRL) offers a powerful framework for developing such an adaptive AI agent [3, 4], leveraging its ability to learn optimal policies through interactions with both the human partner and the environment [5, 6]. However, employing DRL in HAIC presents two interdependent challenging goals in balancing exploration and exploitation:
@@
i) Enhancing exploration of critical but rare state-action pairs yielding sparse rewards. Rewards are usually sparse in HAIC due to the inherent complexity of coordinated tasks. The state-action pairs associated with sparse rewards in HAIC are critical for effective learning. However, these pairs are rarely encountered during training, since they emerge from coordinated sequences of temporally-extended actions between AI and human agents [7]–making them particularly elusive when the AI agent has limited ability to fully interpret the intricate dynamics of unpredictable human behaviors [8, 9]. These rare interactions represent key moments when AI actions perfectly align with human intentions, leading to successful task completion. Without sufficient exposure to these critical state-action pairs, the AI agent struggles to understand which actions are beneficial, resulting in slow learning and local optima performance. Existing approaches mitigate this issue by augmenting exploration through intrinsic rewards that supplement those sparse rewards obtained extrinsically from the environment [10, 11].
Figure 1: The proposed BCR-DRL for HAIC. The AI agent is trained by our BCR-DRL algorithm, which integrates extrinsic rewards, intrinsic rewards, and context-aware weights. Extrinsic rewards are obtained from the environment, whilst intrinsic rewards are developed based on the behaviors of both human and AI agents. Context-aware weights are dynamically adjusted according to the training context across three domains, such as task performance, AI agent behaviors, and human behaviors.
@@
ii) Ensuring effective exploitation of the explored state-action pairs. Although the above sparse rewarding challenge in HAIC can be mitigated by intrinsic rewards, continuously relying on them throughout training can lead to unstable performance [12, 13, 14], due to the inherent exploration-exploitation trade-off in DRL [15, 16, 17]. To address this issue, we can leverage the training context in HAIC to optimize the exploration and exploitation. Specifically, the context in HAIC captures the evolving state of coordination between AI and human participants, such as accumulated sparse rewards, rewards instability conditions, and learning progress. This contextual information enables AI agents to adaptively adjust their learning strategy: prioritizing exploration in early stages when critical state-action pairs are scarce, while gradually transitioning to exploitation in later stages to refine the learned coordination patterns.
@@
In this paper, we design an innovative behavior- and context-aware reward (BCR) for DRL, namely BCR-DRL, to address the aforementioned challenges in HAIC. Our BCR (see Fig. 1) extends conventional extrinsic rewards with two key components: (i) Dual intrinsic rewards enhance exploration by encouraging both diverse AI actions and distinct behavior from the human partner, and (ii) Context-aware weights that optimize exploitation by dynamically adjusting reward weights based on the training context. Specifically, the first component includes two intrinsic rewards: AI self-motivated rewards that promote action diversity, and human-aware rewards that help understand human intentions through counterfactual reasoning. Both rewards utilize logarithmic terms to increase the likelihood of encountering critical but rare state-action pairs. The second component dynamically balances intrinsic and extrinsic rewards based on coordination effectiveness, measured through task completion rates, human behavior patterns, and agent diversity. Experimental results demonstrate that our BCR-DRL outperforms state-of-the-art algorithms, achieving higher sparse rewards and faster training convergence, which illustrate an innovative path for effective and seamless HAIC.
@@ 2 Related Work
Intrinsic rewards. Prior work in multi-agent reinforcement learning has demonstrated the effectiveness of social influence as intrinsic rewards [10]. This approach uses counterfactual actions of the ego AI agent to encourage behaviors that significantly influence other AI agents’ actions. However, HAIC presents fundamentally different challenges compared to standard MARL scenarios. Unlike MARL where all agents are trainable via RL algorithms, HAIC involves uncontrollable humans whose policies cannot be optimized through RL. This fundamental difference necessitates human behavior modeling approaches (such as behavior cloning) that are incompatible with standard MARL frameworks like centralized training with decentralized execution (CTDE) [18]. More recent work has proposed a model-free approach that maximizes AI behavior diversity without explicitly modeling human behavior [5]. While this approach enables adaptation to general human behaviors, the no free lunch theorems [19] suggest that optimal performance in specific scenarios requires exploiting knowledge about the particular behavioral patterns of individuals or groups.
We bridge this gap by proposing a dual intrinsic rewarding scheme that combines the benefits of both approaches while addressing their limitations. Our approach composes: (i) an AI self-motivated intrinsic reward that facilitates comprehensive exploration through behavioral diversity, and (ii) a human-motivated intrinsic reward that utilizes counterfactual human actions obtained from the pre-trained human model, enabling the AI agent to better adapt human intentions. This dual approach enables the AI agent to maintain a comprehensive exploration of its action space while developing actions that effectively complement specific human behavioral patterns, addressing the fundamental constraint that human policies cannot be directly optimized in HAIC settings.
Critical rare state-action pairs. Capturing the critical state-action pairs associated with the sparse rewards is crucial for efficient DRL training in complicated scenarios like HAIC. However, these pairs are usually rarely encountered since they are the results of executing a series of temporally-extended actions [7, 8, 9]. Without sufficient exposure of these critical rare state-action pairs, the AI agent struggles to identify effective policies, resulting in slower learning and local optima performance.
@@
Adaptive weights. Although incorporating intrinsic rewards can encourage the exploration of the AI agent, it introduces the risk of over-exploration and low data sample utilization efficiency. This issue becomes particularly acute when critical state-action pairs, which were initially rare, but being captured more frequently as training progresses. At this point, continuing to emphasize exploration through intrinsic rewards may hinder rather than help the learning process. Therefore, a desirable DRL reward should be context-aware–capable of recognizing and adjusting when to emphasize the exploration by enlarging the intrinsic rewards and when to emphasize the exploitation by strengthening the extrinsic rewards.
@@
The key challenge here lies in dynamically balancing these competing objectives based on the current training context. This bears a similarity to balancing multiple tasks in deep learning [20]. For example, adaptive weighting has been developed to dynamically adjust task weights based on the instabilities caused by the co-existing tasks in deep learning [21]. Inspired by this concept of adaptive weighting, we propose a novel context-aware weighting mechanism for HAIC. Our approach dynamically adjusts the weights between different rewards by monitoring the training context, including the accumulated sparse rewards and instabilities caused by different rewards, aiming at efficiently achieving the global optimal of sparse rewards.
3 Preliminaries
3.1 Foundations for Logarithmic Intrinsic Rewards
We provide the theoretical foundations to support the use of logarithmic forms in intrinsic reward design, which will be detailed in eqs. (6) and (7). Intuitively, this design focuses on log-likelihood, increasing the sensitivity of the policy to rate state-action pairs.
Theorem 1 (Entropy-like Logarithmic Intrinsic Rewards for Rare State-action Pair Oriented Policy Updates).
Let $\pi_{\theta}(a|o)$ be a parameterized stochastic policy over action $a$ given observation $o$. For the counterfactual scenario, let $\tilde{o}$ be a counterfactual observation (e.g., conditioned on human-only action), and $\tilde{a}$ represents the counterfactual action taken with the same policy but under the counterfactual observation. The unified logarithmic intrinsic reward is defined as
$\mathcal{R}(a,o,\tilde{o}|\pi_{\theta},\delta)\!\!=\!\!\left|\log(\pi_{\theta}(a|o))\!-\!\log(\pi_{\theta}(\tilde{a}|\tilde{o}))^{\delta}\right|,\delta\in\{0,1\},$
where $\delta$ determines the specific form of the reward: $\delta=0$ reduces to $|\log\pi_{\theta}(a|o)|$, and $\delta=1$ yields a symmetric difference between log-likelihoods. As a result, when optimizing the expected intrinsic reward $\mathbb{E}_{\pi_{\theta}}[\mathcal{R}(a,o,\tilde{o}|\pi_{\theta},\delta)]$, the resulting policy gradients naturally amplify updates toward rare state-action pairs, enhancing adaptability in environments with sparse external rewards.
Proof.
We prove that the logarithmic intrinsic reward amplifies policy updates toward rare state-action pairs by analyzing the reward scaling properties.
For the case $\delta=0$, the reward reduces to $\mathcal{R}(a,o|\pi_{\theta})=|\log\pi_{\theta}(a|o)|$. Since $0<\pi_{\theta}(a|o)\leq 1$, we have $\log\pi_{\theta}(a|o)\leq 0$, thus $|\log\pi_{\theta}(a|o)|=-\log\pi_{\theta}(a|o)$.
Consider two actions with probabilities $p_{1}>p_{2}>0$. The ratio of their logarithmic intrinsic rewards is:
$\frac{\mathcal{R}(a_{2})}{\mathcal{R}(a_{1})}=\frac{-\log p_{2}}{-\log p_{1}}=\frac{\log(1/p_{2})}{\log(1/p_{1})}$ (1)
Since $p_{1}>p_{2}$, we have $1/p_{2}>1/p_{1}$, therefore $\frac{\mathcal{R}(a_{2})}{\mathcal{R}(a_{1})}>1$, showing that rarer actions receive higher intrinsic rewards.
More importantly, if $p_{2}=\epsilon\cdot p_{1}$ where $0<\epsilon<1$, then:
$\frac{\mathcal{R}(a_{2})}{\mathcal{R}(a_{1})}=\frac{\log(1/(\epsilon p_{1}))}{\log(1/p_{1})}=1+\frac{\log(1/\epsilon)}{\log(1/p_{1})}$ (2)
As $\epsilon\to 0$ (action becomes rarer), this ratio grows unboundedly, demonstrating exponential amplification for rare actions.
In contrast, traditional entropy-based rewards $\mathcal{H}(a|o)=-\pi_{\theta}(a|o)\log\pi_{\theta}(a|o)$ yield:
$\frac{\mathcal{H}(a_{2})}{\mathcal{H}(a_{1})}=\frac{p_{2}\log(1/p_{2})}{p_{1}\log(1/p_{1})}=\epsilon\cdot\frac{\log(1/(\epsilon p_{1}))}{\log(1/p_{1})}$ (3)
As $\epsilon\to 0$, this ratio approaches 0, showing diminishing rewards for rare actions.
The policy gradient contribution is proportional to $\mathcal{R}(a,o|\pi_{\theta})\nabla_{\theta}\log\pi_{\theta}(a|o)$. For logarithmic rewards, rare actions (small $p$) receive gradient amplification of $O(-\log p)$, while entropy-based methods provide only $O(-p\log p)$, which vanishes as $p\to 0$.
Therefore, the logarithmic form naturally amplifies updates toward rare state-action pairs, enhancing exploration in sparse reward environments. This completes the proof. ∎
Example 1.
Consider a stochastic policy $\pi_{\theta}(a|o)$ over 333 actions in a given observation state $o_{1}$, with probabilities $\pi_{\theta}(a_{1}|o_{1})=0.7$, $\pi_{\theta}(a_{2}|o_{1})=0.2$, and $\pi_{\theta}(a_{3}|o_{1})=0.1$. We compare the traditional entropy-based intrinsic reward $\mathcal{H}(a_{i}|o_{i})=-\pi_{\theta}(a_{i}|o_{i})\!\log(\pi_{\theta}(a_{i}|o_{i})),\forall i\in\{1,2,3\}$ with our proposed logarithmic intrinsic reward $\mathcal{R}(a_{i},o_{i}|\pi_{\theta})=\left|\log(\pi_{\theta}(a_{i}|o_{i}))\right|,\forall i\in\{1,2,3\}$ (corresponding to $\delta=0$ in Theorem 1):
For the traditional entropy-based form, we have
$\mathcal{H}(a_{1}|o_{1})=-\pi_{\theta}(a_{1}|o_{1})\log(\pi_{\theta}(a_{1}|o_{1}))=0.250$,
$\mathcal{H}(a_{2}|o_{1})=-\pi_{\theta}(a_{2}|o_{1})\log(\pi_{\theta}(a_{2}|o_{1}))=0.322$,
$\mathcal{H}(a_{1}|o_{1})=-\pi_{\theta}(a_{3}|o_{1})\log(\pi_{\theta}(a_{3}|o_{1}))=0.230$.
For our logarithmic form, we have
$\mathcal{R}(a_{1}|o_{1})=\left|\log(\pi_{\theta}(a_{1}|o_{1}))\right|=\left|\log(0.7)\right|=0.357$,
$\mathcal{R}(a_{2}|o_{1})=\left|\log(\pi_{\theta}(a_{2}|o_{1}))\right|=\left|\log(0.2)\right|=1.609$,
$\mathcal{R}(a_{3}|o_{1})=\left|\log(\pi_{\theta}(a_{3}|o_{1}))\right|=\left|\log(0.1)\right|=2.303$.
Remark 1.
In Example 1, the traditional entropy-based form, $\mathcal{H}(\cdot)$, provides relatively balanced rewards across all actions, with the highest reward value of 0.3220.3220.322 for action $a_{2}$. In contrast, our logarithmic intrinsic reward form, $\mathcal{R}(\cdot)$, particularly amplifies the reward for rare actions, assigning 2.3032.3032.303 to the least likely action $a_{3}$, a value approximately 6.56.56.5 times higher than the reward for the most common action $a_{1}$. Compared with the traditional entropy-based method, this property creates a stronger incentive for policy updates toward exploring rare state-action pairs, which is particularly valuable when critical behaviors might be associated with low-probability actions. By removing the probability weighting term, the logarithmic form establishes a reward landscape that more effectively promotes exploration of the entire state-action space, thereby facilitating the HAIC scenarios that tend to experience sparse rewards.
We also note that the logarithmic form offers computational simplicity by eliminating the probability multiplier term, reducing computational complexity compared to entropy-based methods while maintaining effective exploration capabilities.
3.2 Benchmark Human-AI Coordination
@@
Experimental coordination aspects. We first evaluate our approach using Overcooked [1, 22], a standardized benchmark for HAIC that simulates collaborative cooking tasks requiring coordinated actions between human and AI agents. This environment is particularly suitable for studying HAIC, emphasizing sparse rewards, as successful task completion requires executing precise sequences of coordinated actions of both human and AI agents. We conduct experiments across three distinct layouts of the overcooked environment. Each of the three layouts highlights different aspects of coordination between human and AI agents. The Cramped Room layout focuses on coordination between agents located in a shared space, where collision avoidance between human and AI agents is critical; The Asymmetric Advantages layout concentrates on coordination between agents located in distinct areas with varying access to cooking resources, where asymmetric behavior planning is critical; The Coordination Ring layout evaluates coordination between agents located in a small room with a central obstacle, where path planning is critical.
To further investigate the generalization ability of our approach beyond Overcooked, we introduce an Exploration environment, where a human and an AI agent must coordinate to explore a shared space by jointly covering all accessible areas within a limited number of timesteps. This setting focuses on coordination in terms of asynchronous and complementary actions, where agents must avoid redundant movements, adapt to each other’s partial progress, and efficiently divide the exploration workload. Detailed analysis along with the experiments is given in Section 5.
@@
Human behavior alignment. Current DRL-based AI agents for HAIC fall into two categories: model-free approaches that train AI agents to adapt to general human behaviors without relying on specific human model [5], and model-based approaches that enable more personalized adaptation to individual or group-specific behavioral patterns [1, 6, 23]. Our work adopts the model-based approach. The human models utilized in the Overcooked layouts are developed by the authors of [1]. These real human-cloned models use data collected from a group of humans to imitate the behaviors of human players across three different layouts in the Overcooked environment. The reliability of the selected human models has been validated by the original Overcooked creators [1]. Subsequent studies [5, 6] further confirm their ability to reproduce human-level coordination patterns in all three layouts. For the generalization study, different from the human-cloned models used in Overcooked, the human model is a deliberately synthetic agent designed to capture more generic and stochastic behaviors, focusing on coordination under temporal asynchrony and behavioral uncertainty. Details of this setup are provided in Section 5.3.
Remark 2.
@@
The selection of the model-based approach ensures reliable validation of our BCR-DRL algorithm, particularly its human-motivated intrinsic reward component, which builds upon accurate modeling of human behavior. The three layouts are selected based on their demonstrated reliability in previous work where trained human models achieved performance comparable to average human players [1]. While our BCR-DRL algorithm’s performance depends on human model accuracy, it remains robust as long as behavioral patterns are consistent between training and testing phases, regardless of the specific way the human model behaves. This robustness was demonstrated in Section 5.3, where we achieved performance improvements even with a simplified random behavior model that differs significantly from real human behavior patterns.